OpenStudy (anonymous):

Solve for x: 2x – 8 = x – 9

OpenStudy (anonymous):

-1

terenzreignz (terenzreignz):

ok, step by step, but try to do the next ones by yourself, ok? 2x - 8 = x - 9 the x on the right side, we bring it to the left side, but change the sign, so..... 2x - 8 - x = - 9 the (-8) on the left side, we bring it to the right side, but change the sign, so.... 2x - x = - 9 + 8 Then you just simplify to... x = -1 OK? :)